OpenStudy (anonymous):

Factor completely 2x^2-x-6

OpenStudy (anonymous):

@Vlery

OpenStudy (vlery):

(2x+3)(x-2)

OpenStudy (anonymous):

\[2x ^{2}-x-6=2x ^{2}+2x-3x-6=2x(x+1)-3(x+2)=(x+2)(2x-3)\]